You are standing on the Earth at a location on the equator where your distance from the center of the Earth is r = 6.53×106 m. What is your linear velocity due to the Earth's rotation?
Answer #1

Solution:

Given:

R = 6.53 x 106 m

Since: angular velocity of earth () = 2 / (24 x 60 x 60) rad/s = 7.27 x 10-5 rad/s

Therefore: linear velocity = R = ( 6.53 x 106 m)(7.27 x 10-5 rad/s) = 474.78 m/s

Thus: your linear velocity due to the Earth's rotation is : 474.78 m/s